Shift report 20030118 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Jan 18 2003 - 00:16:28 EST

  • Next message: Apache: "Shift report 20030118 00:00-08:00"
    ------------------------------------------------------------
    Date: Sat Jan 18 00:16:27 2003, Shift: 16:00-24:00
    Supervisor: JJG, with: 
    Log book#13 pages 79 - 83
    
    Shift Summary:
    ------------------------------------------------------------
    1. beam since 21.15. High background initially . H1/inel =30
       background decreased to H1/inel= 1 at 22.50. Told MCR to
       keep settings as role model.
    2. Many crashes of HV, crates, DAQ etc. Finally start data 
       taking at 22.50 run 7444.
    3. runs 7440-7443 are useless since various Hv are off.
    4. FS trigger, H2 ch 31 fixed. 
    5. H1 first 16 slats still weird. software?? (pedestal
       sub.??)
    6. Width of TOFW(or) increased to 40 ns. Precedes INEL and
       MRS-T0 by about 30 ns. 
    7. All anodes were on after last shift with no beam: 
       Shifters: REMEMBER TO TURN DOWN GAS DET. ANODES when 
       beam disappears!! 
    8. We ought to use ONLY E-LOG soon. Redundancy with 2 log
       books is tedious.
    9. Need international tel. in CH to be able to call   
       collaborators looking at data in Europe.  
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat Jan 18 2003 - 14:56:27 EST